# VEDA: 3D Molecular Generation via Variance-Exploding Diffusion with Annealing

### v1

@@ Conclusions
In this work, we introduced VEDA, a novel equivariant generative model for 3D molecular structures that unifies continuous and discrete generative processes within a single framework. Through comprehensive experiments on QM9 and GEOM-DRUGS, we show that VEDA not only matches or surpasses strong baselines like UniGEM and SemlaFlow in atom- and molecule-level stability, but also delivers efficient generation thanks to its unified diffusion framework. Ablations confirm that both the amplified VE noise schedule and the discrete sampling mechanism are critical to VEDA’s success.
The success of VEDA in unconditional generation establishes a robust foundation, opening a clear and immediate path toward property-guided molecular design. Building on the smooth energy landscapes VEDA creates, our direct next step is to condition the generative process on key chemical properties, such as binding affinity or solubility. On a more fundamental level, we believe a key limitation of the current approach—common in the field—is its reliance on implicit velocity prediction. Future equivariant architectures should be designed to explicitly output the time-dependent vector field $v_{\theta}(\mathbf{x},t)$. This would align training objectives with advanced samplers like Consistency Flow Matching (lusimplifying) and MeanFlow (geng2025mean), further improving sampling efficiency and enabling one- or few-step integration.
@@ Appendix C Ablation Studies
Appendix D Molecule Algorithm
@@
Appendix E Experiments of various sampling steps
@@
Appendix F Derivation of Transition Rates Matrix for the Masking Process
@@
We need to make some assumptions about the form of $R^{\mathrm{DB}}$. When considering masking noise, a process that is in detailed balance will have a certain rate for transitions from a masked state to $x_{0}$, as well as a certain rate for transitions from $x_{0}$ back to the masked state in order to cancel out this effect. The rate for such transitions would take the following form. So We have parameterized $R^{\mathrm{DB}}$ with two time-dependent pre-defined constants $a_{t}$ and $b_{t}$. Therefore, when $i$ and $j$ are both not equal to $x_{0}$, this term is zero, essentially representing the diffusion flow between $x_{0}$ and other states and its reverse action.
$R^{\mathrm{DB}}(i,j)=a_{t}\,\delta(i,x_{0})+b_{t}\,\delta(j,x_{0}).$ (21)
$p(i)=(1-m(t))\,\delta(i,x_{0})+\frac{m(t)}{S}$ (22)
@@
$\bigl((1-m)\,\delta(i,x_{0})+\tfrac{m}{S}\bigr)\bigl(a_{t}\,\delta(i,x_{0})+b_{t}\,\delta(j,x_{0})\bigr)$ (23)
$=$ $\bigl((1-m)\,\delta(j,x_{0})+\tfrac{m}{S}\bigr)\bigl(a_{t}\,\delta(j,x_{0})+b_{t}\,\delta(i,x_{0})\bigr).$ (24)
For the case $x_{0}=i$, $\delta(i,x_{0})=1$ and $\delta(j,x_{0})=0$, we have:
@@
$R^{\mathrm{DB}}(i,j)=\eta\,\delta(i,x_{0})+\eta\,\frac{S-S\,m(t)+m(t)}{m(t)}\,\delta(j,x_{0}).$ (27)
@@
$R^{\theta}(x_{t},j)$ $=\mathbb{E}_{p^{\theta}_{0|t}(x_{0}\mid x_{t})}\left[R^{*}(x_{t},j)+R^{\mathrm{DB}}(x_{t},j)\right]$ (28)
$=\mathbb{E}_{p^{\theta}_{0|t}(x_{0}\mid x_{t})}\left[\frac{m^{\prime}(t)}{m(t)}\,\delta(j,x_{0})\bigl(1-\delta(x_{t},x_{0})\bigr)\right$.
$\qquad\left.+\eta\,\delta(x_{t},x_{0})+\eta\,\frac{S-S\,m(t)+m(t)}{m(t)}\,\delta(j,x_{0})\right]$ (29)
$=\frac{\eta S(1-m(t))+\eta m(t)+m^{\prime}(t)}{m(t)}p^{\theta}_{0|t}(x_{0}=j\mid x_{t})$
$\qquad+\eta p^{\theta}_{0|t}(x_{0}=x_{t}\mid x_{t})$ (30)
In this formulation, the term $p^{\theta}_{0|t}(x_{0}=j\mid x_{t})$ represents the predicted probability that the clean data corresponds to category $j$, and thus determines the transition rate from the current state $x_{t}$ to $j$. The second term, $p^{\theta}_{0|t}(x_{0}=x_{t}\mid x_{t})$, represents the probability that the current noisy state $x_{t}$ already matches the clean data, and therefore governs the rate of remaining in the current state.
@@ Appendix G Derivation of Transition Rates Matrix for the Masking Process
Appendix H Derivation of Optimal $\alpha_{t}$
@@
Appendix I Derivation for the Smoothed Potential Energy Equivalence

### v2

@@ Conclusions
In this work, we introduced VEDA, a novel equivariant generative model for 3D molecular structures that unifies continuous and discrete generative processes within a single framework. Through comprehensive experiments on QM9 and GEOM-DRUGS, we show that VEDA not only surpasses strong baselines like GeoBFN and SemlaFlow in structural validity and stability metrics, but also delivers efficient generation thanks to its unified diffusion framework. Ablations studies validate that both the amplified VE noise schedule and the discrete sampling mechanism are critical to VEDA’s performance.
The success of VEDA in unconditional generation establishes a robust foundation, opening a clear and immediate path toward property-guided molecular design. A natural extension is to condition the generative process on target chemical properties such as binding affinity or solubility, leveraging VEDA’s unified framework to guide generation toward desired molecular profiles. Beyond conditional generation, we identify a fundamental limitation shared across current diffusion models: reliance on implicit velocity prediction through score matching. Future equivariant architectures should be designed to explicitly output the time-dependent vector field $v_{\theta}(\mathbf{x},t)$. This would align training objectives with advanced samplers like Consistency Flow Matching (lusimplifying) and MeanFlow (geng2025mean), further improving sampling efficiency and unlocking one- or few-step integration.
@@ Appendix A Experiment Details
Sampling Temperature ($\tau$) 1.0 0.9 1.0 1.0
Max $w_{\text{cat}}$ (Clip) N/A 10 N/A 10
Prediction Mode Constant Constant Constant Adaptive
Cat. Loss Weighting N/A $1/m(t)$ N/A $1/m(t)$
@@
Self-Conditioning
Following the configuration of the SemlaFlow baseline (irwin2025semlaflow), we incorporate a self-conditioning mechanism for the VEDA-S model to enhance generation quality. This technique involves augmenting the model input with a preliminary estimate of the clean data, denoted as $\tilde{x}_{0}$.
  • Training: We employ a stochastic conditioning strategy. With a probability of $0.5$, the model performs an initial forward pass to estimate $\tilde{x}_{0}$, which is then concatenated with the noisy input for the final prediction. In the remaining $50\%$ of cases, the conditioning input is set to zero (null condition).
  • Sampling: Self-conditioning is active for all inference steps except the initial one (where no prior estimate exists). The estimate of the clean data derived from the previous timestep $t_{i-1}$ is used as the condition for the current step $t_{i}$.
Note that this mechanism is exclusive to VEDA-S and is not employed in the VEDA-E architecture.
@@ Appendix C Ablation Studies
Appendix D Additional Implementation Details
This section elaborates on the hyperparameter configurations, stabilization mechanisms, and architectural variants used in our experiments.
Alternative Model Configurations
To evaluate the robustness of our method, we provide comparisons between different residual prediction modes and mask rate schedules:
  • Residual Prediction Mode: This configuration dictates how the input is subtracted from the network output to formulate the denoising target.
    • Adaptive (Default): Uses the optimal LMMSE coefficient derived in Appendix I. The network predicts a residual where the input is scaled by a time-dependent factor $\alpha_{t}$:
      $\mathbf{x}_{\text{out}}=F_{\theta}(\mathbf{x}_{\text{in}})-\alpha_{t}\mathbf{x}_{\text{in}},\quad\text{where }\alpha_{t}=\frac{\sigma_{d}t}{\sigma_{d}^{2}+t^{2}}.$
    • Constant: A simplified residual connection where the input is subtracted directly without scaling (effectively setting $\alpha_{t}=1$). This forces the network to predict the pure residual $\mathbf{x}_{\text{out}}=F_{\theta}(\mathbf{x}_{\text{in}})-\mathbf{x}_{\text{in}}$.
  • Mask Rate Schedule: This determines the mapping from the continuous noise level $t$ to the discrete mask rate $m(t)\in[0,1]$.
    • Log-uniform (Default): Linearly maps the logarithmic noise level $\ln(t)$ to the mask rate $m(t)$. This ensures that the discrete masking process is uniformly distributed with respect to the logarithmic time scale used by the continuous diffusion.
    • Sigmoid-style (EDM): Adopts a sigmoid-like mapping derived from the ratio of noise to data weight, commonly used in discrete flow matching to map unbounded diffusion time $t\in[0,\infty)$ to a bounded probability:
      $m(t)=\frac{t}{t+\sigma_{d}}.$
Training and Sampling Dynamics
We incorporate two specific mechanisms to stabilize training and refine the generation process:
  • Categorical Loss Weighting: While heuristic objectives like MaskGIT (chang2022maskgit) have shown promise, they often lack the scaling terms required for a rigorous link to maximum likelihood estimation (nie2025large). To bridge this gap and balance the training objective, we employ a time-dependent weighting scheme inversely proportional to the mask rate:
    $w_{\text{cat}}(t)=\min\left(\frac{1}{m(t)},10\right).$
    By incorporating this $1/m(t)$ factor, we ensure that time steps with fewer masked tokens (low $m(t)$) are not under-represented in the total loss. The clipping mechanism further prevents numerical instability near $t=0$, effectively combining theoretical motivation with training stability.
  • Sampling Temperature: During the generation phase, we utilize a temperature parameter $\tau$ to rescale the logits of the categorical distributions (atom types, charges, and bond types) prior to the Softmax operation: $P(z_{i})=\text{Softmax}(l_{i}/\tau)$, where $l_{i}$ are the predicted logits. This parameter controls the stochasticity of the discrete transitions, balancing diversity with structural validity.
Appendix E Molecule Algorithm
@@
Appendix F Experiments of various sampling steps
@@ Appendix G Derivation of Transition Rates Matrix for the Masking Process
We need to make some assumptions about the form of $R^{\mathrm{DB}}$. When considering masking noise, a process that is in detailed balance will have a certain rate for transitions from a masked state to $z_{0}$, as well as a certain rate for transitions from $z_{0}$ back to the masked state in order to cancel out this effect. The rate for such transitions would take the following form. So We have parameterized $R^{\mathrm{DB}}$ with two time-dependent pre-defined constants $a_{t}$ and $b_{t}$. Therefore, when $i$ and $j$ are both not equal to $z_{0}$, this term is zero, essentially representing the diffusion flow between $z_{0}$ and other states and its reverse action.
$R^{\mathrm{DB}}(i,j)=a_{t}\,\delta(i,z_{0})+b_{t}\,\delta(j,z_{0}).$ (25)
$p(i)=(1-m(t))\,\delta(i,z_{0})+\frac{m(t)}{S}$ (26)
@@
$\bigl((1-m)\,\delta(i,z_{0})+\tfrac{m}{S}\bigr)\bigl(a_{t}\,\delta(i,z_{0})+b_{t}\,\delta(j,x_{0})\bigr)$ (27)
$=$ $\bigl((1-m)\,\delta(j,z_{0})+\tfrac{m}{S}\bigr)\bigl(a_{t}\,\delta(j,z_{0})+b_{t}\,\delta(i,z_{0})\bigr).$ (28)
For the case $x_{0}=i$, $\delta(i,z_{0})=1$ and $\delta(j,z_{0})=0$, we have:
@@
$R^{\mathrm{DB}}(i,j)=\eta\,\delta(i,z_{0})+\eta\,\frac{S-S\,m(t)+m(t)}{m(t)}\,\delta(j,x_{0}).$ (31)
@@
$R^{\theta}(z_{t},j)$ $=\mathbb{E}_{p^{\theta}_{0|t}(z_{0}\mid z_{t})}\left[R^{*}(z_{t},j)+R^{\mathrm{DB}}(z_{t},j)\right]$ (32)
$=\mathbb{E}_{p^{\theta}_{0|t}(z_{0}\mid z_{t})}\left[\frac{m^{\prime}(t)}{m(t)}\,\delta(j,z_{0})\bigl(1-\delta(z_{t},z_{0})\bigr)\right$.
$\qquad\left.+\eta\,\delta(z_{t},z_{0})+\eta\,\frac{S-S\,m(t)+m(t)}{m(t)}\,\delta(j,z_{0})\right]$ (33)
$=\frac{\eta S(1-m(t))+\eta m(t)+m^{\prime}(t)}{m(t)}p^{\theta}_{0|t}(z_{0}=j\mid z_{t})$
$\qquad+\eta p^{\theta}_{0|t}(z_{0}=z_{t}\mid z_{t})$ (34)
In this formulation, the term $p^{\theta}_{0|t}(z_{0}=j\mid z_{t})$ represents the predicted probability that the clean data corresponds to category $j$, and thus determines the transition rate from the current state $z_{t}$ to $j$. The second term, $p^{\theta}_{0|t}(z_{0}=z_{t}\mid z_{t})$, represents the probability that the current noisy state $z_{t}$ already matches the clean data, and therefore governs the rate of remaining in the current state.
@@
Appendix H Derivation of Transition Rates Matrix for the Masking Process
@@
Appendix I Derivation of Optimal $\alpha_{t}$
@@
Appendix J Derivation for the Smoothed Potential Energy Equivalence